Approximately how many 20cm x 20cm square tiles are needed to cover a square floor whose side is 16m?
1
Expert's answer
2012-06-12T10:47:58-0400
S1=20*20=400 cm2=0.04 m2
S2=16 m*16 m=256 m2
N=S2/S1=6400
Answer: N=6400 tiles